OpenStudy (anonymous):

i need help with these problems 2. 3x + 2y = 7 –3x + 11 = y (1 point) (6, –3) (6, –7) 5. 6x + 2 = y 3y – 18x = 12 one solution two solutions infinitely many solutions no solution

OpenStudy (anonymous):

your first question (5,-4) solves----- the two shown don't ----3*5+2*(-4)=15-8=7 y=6 x+2 so 3 y=18 x+6 first equation 3 y=18 x+12 second equation these can never be equal